OpenStudy (luigi0210):

Limit Question:

OpenStudy (luigi0210):

\[\lim_{x \rightarrow 0} \frac{x}{\sin3x}\]

OpenStudy (amistre64):

you know a Lhopital?

OpenStudy (luigi0210):

Yes, but I don't know what to do with that sin3x

OpenStudy (amistre64):

whats the derivative of sin(3x)?

OpenStudy (yrelhan4):

Lhospital will work. Or you should know the identity lim y-->0 sin(y)/y and y/siny is 1.. In this question y=3x.

OpenStudy (amistre64):

\[sin(u)=u-\frac{1}{3!}u^3+\frac{1}{5!}u^5\pm...\] \[sin(3x)=3x-\frac{1}{3!}(3x)^3+\frac{1}{5!}(3x)^5\pm...\] \[\frac1xsin(3x)=3-\frac{1}{3!}(3x)^2+\frac{1}{5!}(3x)^4\pm...\] at x=0, that gives us 3 \[\lim\frac{1}{\frac1xsin(3x)}=\frac13\]

OpenStudy (amistre64):

sin(3x) derives to 3 cos(3x) sooo x/sin(x) to 1/3cos(3x); at x=0 is 1/3(1) = 1/3 as well

OpenStudy (luigi0210):

Alright, thanks :)

OpenStudy (amistre64):

good luck